Consider the following balanced reaction. What mass (in g) ofCO2 can be formed from 288 mg of O2? Assumethat there is excess C3H7SH present.

C3H7SH(l) + 6 O2(g) → 3CO2(g) + SO2(g) + 4 H2O (g)
Answer #1
288 mg O2 *(1 g/1000mg) *(1mol /32 g O2) *(3 mol CO2 / 6 mol O2)*(44 g/1mol CO2) = 0.198 grams CO2
Answer #2
        The balencedchemical reaction is
                                                     C3H7SH (l) + 6 O2 (g) -------> 3 CO2 (g) + SO2 (g) +4 H2O (g)
                               
       From above reaction whatwe can understand that 6.0 mols of O2   gives 3.0 mols ofCO2 .
         So now 288 mgof O2 formed CO2 is    ------g
               6 mols of O2 means 6*32g   = 192 g
             288 mg of O2 means   288*10-3 g = 0.288g
         1mol ofCO2 = 44g of CO2
  
                             Theformation of CO2 from 288mg of O2   = [(3*44)/ 6*32] *288*10-3 g
                                                                                               = (132/192)*0.288 g
                                                                                               = 0.6875*0.288 g
                                                                                               = 0.198 g
                ∴The mass   of CO2 is formed is 0.198g